3 Consider a firm with an EBIT of $850,000. The firm finances its assets with $2.500,000 debt (costing 7.5 percent) and 400,000 shares of stock selling at $5.00 per share. To reduce the firms risk associated with this financial leverage, the firm is considering reducing its debt by $1,000,000 by selling an additional 200,000 shares of stock. The firm is in the 40 percent tax bracket The change in capital structure will have no effect on the operations of the firm. Thus, EBIT will remain at $850,000. nts Calculate the change in the firms EPS from this change in capital structure. (Round your answers to 2 decimal places.) eBook Ask EPS before EPS amer Print References Diference Lt
Answer #1

Solution

Original EPS:

EBIT=$850000

Interest expense=7.5% of 2500000

=$187500

EBT=850000-187500

$662500

Taxes(40%) =$265000

Net income=662500-265000

=$397500

Shares outstanding=400,000 shares

EPS=$397500/400,000

=$.99375 per share

New EPS after capital structure changes:

Now total number of shares =400000+200000

=600,000

Debt has reduced to =$2,500,000-$1,000,000

=$1, 500,000

New EPS:

EBIT=$850, 000

Interest expense=7.5% of 1,500,000

=$112, 500

EBT=$850, 000-$112, 500

=$737, 500

Taxes 40%=$295, 000

Net income=$442, 500

EPS=$442,500/600,000 shares

=$.7375 per share

EPS is reduced to $ .7375 from $.99375